OpenStudy (anonymous):

If h(x)=(fog)(x) and h(x)=4(x+1)^2, find f(x) and g(x)

OpenStudy (anonymous):

There is no single answer. I mean you could let \[ f(x) = x\\ g(x) = h(x) \]Or \[ f(x) = h(x)\\ g(x) = x \]For example

OpenStudy (anonymous):

In this case though, I think they want you to set \(g(x)\) to the part in parenthesis.

OpenStudy (anonymous):

\[h(x)=\color{blue}{4(}\color{red}{x+1}\color{blue}{)^2}\]

OpenStudy (anonymous):

This would leave you with: \[ g(x) = \color{red}{x+1} \quad\text{and}\quad f(x)=\color{blue}{4x^2} \]

OpenStudy (anonymous):

thanx